# Preferences Evolve And So Should Your Bandits: Bandits with Evolving States for Online Platforms

### v2

@@ 1.2 Related Work
From the online learning literature, our work has connections with papers both on Multi-Armed Bandit (MAB) problems and more general RL settings. There has been a lot of recent interest in settings where the expected rewards of the arms evolve over time (i.e., there is a long-term effect on the system). Levine et al. (2017) and Seznec et al. study “rotting bandits”, where the long-term effect is that as you pull an arm the realized reward presented to the learner decreases. The main difference with our problem is that in “rotting bandits” there is no way to “replenish” what you lost from an arm as you kept pulling it. Additionally, the benchmark policy in rotting bandits is to greedily play the optimal arm at each round, had you known everything in advance, which is not at all the case in our setting.
@@
Kleinberg & Immorlica (2018) study “recharging bandits”, where rewards accrue as time goes by since the last time the arm was played. In “blocking bandits” (Basu et al., 2019, 2021; Bishop et al., 2020) playing an arm makes it unavailable for a fixed number of time slots thereafter. In Heidari et al. (2016); Leqi et al. (2021), the rewards of the arms increase/decrease as they get played. In “rested bandits” (Gittins, 1979) an arm’s expected rewards change only when it is played. In “restless bandits” (Whittle, 1988) rewards evolve independently from the play of each arm. In (Cella & Cesa-Bianchi, 2020) the rewards increase as a function of the time elapsed since the last pull. In “recovering bandits” (Pike-Burke & Grunewalder, 2019) the expected reward of an arm is expressed as a function of the time since the last pull drawn from a Gaussian Process with known kernel. In Warlop et al. (2018), the rewards are linear functions of the recent history of actions. In (Mintz et al., 2020), rewards are a function of a context that evolve according to known deterministic dynamics. In our case, the inherent rewards of the arms do not change; instead, they are filtered through the state which is affected by all previously played arms.
@@
Lykouris et al. (2020) consider the case where the arms have a stochastic component and an adversarial one, which is chosen at each round by adversary. The final mean reward is the product between the stochastic and adversarial components. The difference with our setting is that in our case, stochastic reward is multiplied by the state, which is defined deterministically based on the sequence of prior actions, and cannot be chosen arbitrarily by an adversary. In a similar vein, Gupta et al. (2021) consider the setting where the rewards of pulling different arms are correlated. In our case, rewards are also correlated but they are governed by the state. Correlations also really arise once you pull the arms sequentially, as opposed to their problem, where correlation requires arms to be pulled simultaneously
@@
Our work is also related to RL with MDPs with deterministic transition functions (e.g., (Ortner, 2008; Dekel & Hazan, 2013) for stochastic and adversarial respectively) and with (Ortner & Ryabko, 2012) which studies a stochastic RL setting with a continuous state space. The core difference with our work, however, is that the aforementioned works assume that the learner can observe the state that they find themselves in at each round.
@@ 2.1 Experimental evidence for the functional form in B-DES
If we translate the experiment setup to our model, it would be as if we expose a user to an arm / ad with $(r,b)$ repeatedly. From Eq. (1), the state (inherently tied with the CTR) then changes as:
@@ 3.1 Relaxation: Dynamic Programming with Approximate Rewards
$q_{t+1}\left(H_{1:t}^{\textsf{ALG}}\right)=(1-\lambda)^{t}+\lambda\cdot\sum_{s=0}^{t-1}(1-\lambda)^{t-1-s}\cdot b_{I_{s}}$ (2)
@@
Let $\widehat{\textsf{DP}}$ denote the expected reward of the solution returned by a dynamic programming algorithm with inputs $(\widehat{r}_{i},\widehat{b}_{i})_{i\in[K]}$, where $|\widehat{r}_{i}-r_{i}|\leq\delta$ and $|\widehat{b}_{i}-b_{i}|\leq\delta$. Then, $\widehat{\textsf{DP}}\geq\textsf{OPT}-\delta T$.
@@
Because of the knapsack-style structure, we can create an FPTAS for the problem (see e.g., Williamson & Shmoys (2011)) as in Algorithm 1. To see this, one can round to multiples of $\epsilon$, remove “dominated” tuple components, and end up with at most $t/\epsilon$ points in each $F_{t}$. Formally, a pair $(\rho,q)$ dominates another pair $(\rho^{\prime},q^{\prime})$ if $\rho\geq\rho^{\prime}$ and $q\geq q^{\prime}$.
Input. Estimators $\hat{r}_{i},\hat{b}_{i}$ and approximation parameter $\epsilon$. Tuple sequence initialization: $F_{0}\leftarrow\{(0,1)\}$   // $\rho_{0}=0,q_{0}=1$
@@
       $F_{t}\leftarrow F_{t-1}$   // Start from the previously optimal sequence
@@ 3.2 Estimating the IV Rewards and ES
At a high level, it may seem impossible to disentangle learning the $r_{i}$’s and $b_{i}$’s just by observing the reward realization, which depends on their product. In fact, this hardness persists even if one of the two (either the $r_{i}$’s or the $b_{i}$’s) was known. To see this, we present two simple thought experiments.
@@
Given Lemma 3.5 and Corollary 3.6, we now have a way to disentangle the learning of the $r_{i}$’s and the $b_{i}$’s. This is precisely the idea behind our algorithm: first, build estimators for the $r_{i}$’s and subsequently, use these when trying to infer the $b_{i}$’s. The tricky part arises because of the fact that the $r_{i}$’s and the $b_{i}$’s are connected multiplicatively. Note that for Lemmas 3.73.8 and 3.9 that follow, we use a fixed $\varepsilon$. We tune this $\varepsilon$ optimally in the end to obtain the no-regret guarantee.
Set $\varepsilon,\delta,M$ as stated in Theorem 3.1. Initialize rounds $t=1$. /* Build estimators $\{\widehat{r}_{i}\}_{i\in[K]}$/
@@
2       Initialize reward estimate $\widehat{r}_{i}=0$. for blocks $j\in[M]$ do
@@
4                   Play arm $i_{R}$. Update $t\leftarrow t+1$.
            Play arm $i$, observe reward $R_{j}^{i}$, and update: $\widehat{r}_{i}\leftarrow\widehat{r}_{i}+\nicefrac{{R_{j}^{i}}}{{M}}$.   // $q\approx 1-\varepsilon$
@@
7       Initialize state estimator $\widehat{v}_{i}=0$. for pulls $1,\dots,N(\lambda)$ do
8             Play arm $i$. Update $t\leftarrow t+1$.
@@
             Play arm $i$, observe reward $S_{j}^{i}$, and update: $\widehat{v}_{i}\leftarrow\widehat{v}_{i}+\nicefrac{{S_{j}^{i}}}{{M}}$.   // Play $i$ when $q\approx b_{i}$
10
@@
Play arm $i_{R}$ for $N_{R}$ rounds, updating $t\leftarrow t+1$ after each one.  // State $\geq 1-\varepsilon$
@@
The proofs of Lemmas 3.7 and 3.8 are based on an application of Hoeffding’s inequality combined with Corollary 3.6 to control the time it takes for the state to return to almost $1$.
@@
$\Pr\left[|e_{v}|\geq\frac{\delta}{2}\Big{|}|r_{i}|-|e_{r}|\Big{|}\Big{|}\mathcal{E}_{i}^{\prime}\right]$ $\leq\Pr\left[|e_{v}|\geq\frac{\delta}{2}|r_{i}-\delta|\right]\leq\Pr\left[|e_{v}|\geq\frac{\delta}{2}|r_{i}-\delta|\Big{|}\mathcal{E}_{i}^{\prime}\right]$
$=\Pr\left[|e_{v}|\geq\frac{\delta}{2}|r_{i}-\delta|\right]$
$\leq 2\exp\left(-2M\cdot\left(\frac{\delta}{2}\cdot|r_{i}-\delta|-\varepsilon\right)^{2}\right)$ (Lem. 3.8)
@@
where the third derivation is due to the fact that $\mathcal{E}_{i}^{\prime}$ depends on $e_{r}$ and none of the quantities that we take the conditional on depend on it too. Additionally, the last inequality is due to the fact that $|r_{i}-\delta|\leq 1$. From the law of total probability:
@@
Using Lemmas 3.7 and 3.9 we can prove Theorem 3.1 by bounding the number of rounds it takes for the estimators to converge to approximately correct values with high probability and the regret picked up in the event of failing to converge.
Sketch for the unknown $i_{R}$ case. Note that two instances of arms $\{(r_{i},b_{i})\}_{i\in[K]}$ and $\{(cr_{i},b_{i}/c)\}_{i\in[K]}$ for a scalar $c>0$ are equivalent. So, we can always scale the $b_{i}$’s appropriately to make sure that we have a “replenishing” arm. The next part is to show how Algorithm 2 changes if we do not know which among the $K$ arms is the replenishing arm. The only thing that changes is the way that we estimate $\widehat{r}_{i}$’s; instead of using $i_{R}$ as the benchmark arm, we sample randomly an arm $z$. Then, after enough rounds, we can guarantee that with high probability $\hat{r}_{i}\to\bar{b}r_{i}$, where $\bar{b}=\sum_{j}\nicefrac{{b_{j}}}{{K}}$. The second part of Algorithm 2 remains the same. We can then guarantee that we have obtained estimates $\hat{b}_{i}\to\nicefrac{{b_{i}}}{{\bar{b}}}$ with high probability. Tuning again $\delta,\varepsilon,M$ we obtain the same order regret guarantee. The details and the new algorithm can be found in Appendix B.3.
@@ 4 Slow State Evolution: $\lambda\in[0,\widetilde{\Theta}(1/T)]$
For $\lambda=0$, the problem becomes an instance of the standard stochastic $K$-MAB, since the state is always $q_{t}=q_{0}=1,\forall t$. So, applying the standard UCB algorithm (Auer et al., 2002a) guarantees regret $R_{\textsc{DES}}(T)=\mathcal{O}(\sqrt{/TK\log T})$. For the remainder of the section, we discuss the case where $\lambda\in(0,\widetilde{\Theta}(1/T)]$. Roughly, we show that applying EXP3.P (Auer et al., 2002b) “pretending” that there are no states (i.e., taking $\mathbb{E}[\tilde{r}_{i,t}]$ as being exogenously decided; see Algorithm 6 for formal description) incurs regret that is comparable to $R_{\textsc{DES}}(T)$ up to some factors that we formalize below.
@@
where $I^{*}=\arg\max_{i\in[K]}\sum_{t\in[T]}\tilde{r}_{i,t}=\arg\max_{i\in[K]}\sum_{t\in[T]}\tilde{q}_{t}\cdot r_{i}$. EXP3.P treats the induced sequence of states $\tilde{q}_{t}$ as exogenously given, and hence: $I^{*}=\sum_{t\in[T]}\tilde{q}_{t}\cdot\arg\max_{i\in[K]}r_{i}=\arg\max_{i\in[K]}r_{i}:=i^{\star}$. Using this derivation, Equation (7) becomes:
$R_{\textsc{EXT}}(T)=\sum_{t\in[T]}\tilde{q}_{t}\cdot r_{i^{\star}}-\sum_{t\in[T]}\tilde{r}_{I_{t},t}\leq\mathcal{O}\left(\sqrt{KT\log K}\right)$ (8)
@@
$R_{\textsc{DES}}(T)+\underbrace{\left[\sum_{t\in[T]}\tilde{q}_{t}\cdot r_{i^{\star}}-\sum_{t\in[T]}q_{t}^{\pi^{\star}}r_{\pi^{\star}_{t}}\right]}_{A}\leq\mathcal{O}\left(\sqrt{KT\log K}\right)$
@@
$A$ $=\sum_{t\in[T]}\tilde{q}_{t}\cdot r_{i^{\star}}-\sum_{t\in[T]}q_{t}^{\pi^{\star}}r_{\pi^{\star}_{t}}\geq\sum_{t\in[T]}(1-\lambda)^{t}r_{i^{\star}}-\sum_{t\in[T]}q_{t}^{\pi^{\star}}r_{\pi^{\star}_{t}}{}$ (9)
@@
where the first inequality is because $r_{i^{\star}}\geq r_{i},\forall i\in[K]$ and the second one because $1\geq\hat{q}_{t}$. Since $\{J_{t}\}_{t\in[T]}$ is any sequence, Equation (10) must also hold for the optimal sequence ${\pi^{\star}}$. In other words, $\sum_{t\in[T]}r_{i^{\star}}\geq\sum_{t\in[T]}q_{t}^{\pi^{\star}}r_{\pi^{\star}_{t}}$. Using the latter to relax the RHS of Equation (9) we get:

### v4

@@ 1.2 Related Work
There has also been some work on long-term effects of recommendation systems. For example, johari2018learningabandonment and (yang2024exploration) design learning algorithms for online platforms for recommending content when users are allowed to abandon the platform if they are dissatisfied with the recommendations. maystre2023optimizingaudiorecommendationslongterm design and deploy at industrial scale a podcast recommendation system. Finally, Sumida2023-yw study assortment recommendation algorithms for maximizing the long-term platform revenue when users follow a Multinomial Logit choice model.
@@
From the online learning literature, our work has connections with papers both on Multi-Armed Bandit (MAB) problems and more general RL settings. There has been a lot of recent interest in settings where the expected rewards of the arms evolve over time (i.e., there is a long-term effect on the system). Levine et al. (2017) and Seznec et al. (2020) study “rotting bandits”, where the long-term effect is that as you pull an arm the realized reward presented to the learner decreases. The main difference with our problem is that in “rotting bandits” there is no way to “replenish” what you lost from an arm as you kept pulling it. Additionally, the benchmark policy in rotting bandits is to greedily play the optimal arm at each round, had you known everything in advance, which is not at all the case in our setting. In metelli2022stochasticrisingbandits, the authors study the problem of “Stochastic Rising Bandits”, where the long-term effect is that each arm’s expected reward is affected only by the number of this arm’s pulls, and specifically, the realized reward presented to the learner increases over time. In contrast, in our model each arm’s reward is affected by the all other arms’ pulls and it can both increase and decrease.
@@
Kleinberg and Immorlica (2018) study “recharging bandits”, where rewards accrue as time goes by since the last time the arm was played. In “blocking bandits” (Basu et al., 2019, 2021; Bishop et al., 2020) playing an arm makes it unavailable for a fixed number of time slots thereafter. In Heidari et al. (2016); Leqi et al. (2021), the rewards of the arms increase/decrease as they get played. In “rested bandits” (Gittins, 1979) an arm’s expected rewards change only when it is played. In “restless bandits” (Whittle, 1988) rewards evolve independently from the play of each arm. In (Cella and Cesa-Bianchi, 2020) the rewards increase as a function of the time elapsed since the last pull. In “recovering bandits” (Pike-Burke and Grunewalder, 2019) the expected reward of an arm is expressed as a function of the time since the last pull drawn from a Gaussian Process with known kernel. In Warlop et al. (2018), the rewards are linear functions of the recent history of actions. In (Mintz et al., 2020), rewards are a function of a context that evolve according to known deterministic dynamics. In our case, the inherent rewards of the arms do not change; instead, they are filtered through the state which is affected by all previously played arms.
@@
Lykouris et al. (2020) consider the case where the arms have a stochastic component and an adversarial one, which is chosen at each round by adversary. The final mean reward is the product between the stochastic and adversarial components. The difference with our setting is that in our case, stochastic reward is multiplied by the state, which is defined deterministically based on the sequence of prior actions, and cannot be chosen arbitrarily by an adversary. In a similar vein, Gupta et al. (2021) consider the setting where the rewards of pulling different arms are correlated. In our case, rewards are also correlated but they are governed by the state. Correlations also really arise once you pull the arms sequentially, as opposed to their problem, where correlation requires arms to be pulled simultaneously.
There are some known lower bounds for non-stationary bandit models similar to ours; e.g., 10.5555/2968826.2968849, study the case where all the arms’ rewards are drawn some distributions with mean $\mu_{i}(t)$ changing over time $t$. They prove an $\Omega\left(K^{1/3}V^{1/3}T^{2/3}\right)$ lower bound regret for non-stationary bandits with total variation distance mean rewards upper bounded by $V$.
@@
Our work is also related to RL with MDPs with deterministic transition functions (e.g., (Ortner, 2008; Dekel and Hazan, 2013) for stochastic and adversarial respectively) and with (Ortner and Ryabko, 2012) which studies a stochastic RL setting with a continuous state space. The core difference with our work, however, is that the aforementioned works assume that the learner can observe the state that they find themselves in at each round.
@@ 2.1 Experimental evidence for the functional form in B-DES
If we translate the experiment setup to our model, it would be as if we expose a user to an arm / ad with $(r,b)$ repeatedly. From Eq. (1), the state (inherently tied with the CTR evolution for one arm in the Hohnhold et al. experiment) then changes as:
@@ 3.1 Relaxation: Dynamic Programming with Approximate Rewards
$q_{t+1}\left(H_{1:t}^{\textsf{ALG}}\right)=(1-\lambda)^{t}+\lambda\cdot\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot b_{I_{s}}$ (2)
@@
Let $\widehat{\textsf{DP}}$ denote the expected reward of the solution returned by a dynamic programming algorithm with inputs $(\widehat{r}_{i},\widehat{b}_{i})_{i\in[K]}$, where $|\widehat{r}_{i}-r_{i}|\leq\delta$ and $|\widehat{b}_{i}-b_{i}|\leq\delta$; i.e., $\widehat{\textsf{DP}}\triangleq\max_{i_{1},\dots,i_{T}}\sum_{t\in[T]}[(1-\lambda)^{t}+\lambda\sum_{s\in[t-1]}(1-\lambda)^{t-1-s}\widehat{b}_{i_{t}}]\cdot\widehat{i}_{t}$. Then, $\widehat{\textsf{DP}}\geq\textsf{OPT}-\delta T$.
Lemma 3.4.
Let DP be the true expected reward of the solution returned by a dynamic programming algorithm with inputs $(\widehat{r}_{i},\widehat{b}_{i})_{i\in[K]}$, i.e., $\textsf{DP}\triangleq\sum_{t\in[T]}[(1-\lambda)^{t}+\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}\cdot b_{\widehat{i}^{\star}_{s}}]\cdot r_{\widehat{i}^{\star}_{t}}$. Then, $|\widehat{\textsf{DP}}-\textsf{DP}|\leq\delta T$.
@@
Because of the knapsack-style structure, we can create an FPTAS for the problem (see e.g., Williamson and Shmoys (2011)) as in Algorithm 1. To see this, one can round to multiples of $\epsilon$, remove “dominated” tuple components, and end up with at most $t/\epsilon$ points in each $F_{t}$. Formally, a pair $(\rho,q)$ dominates another pair $(\rho^{\prime},q^{\prime})$ if $\rho\geq\rho^{\prime}$ and $q\geq q^{\prime}$.
Input. Estimators $\hat{r}_{i},\hat{b}_{i}$ and approximation parameter $\epsilon$.
Tuple sequence initialization: $F_{0}\leftarrow\{(0,1)\}$
  // $\rho_{0}=0,q_{0}=1$
@@
Note that as we show in the proof of Theorem 3.1, identifying estimators $(\widehat{r}_{i},\widehat{b}_{i})_{i\in[K]}$ such that $|\widehat{r}_{i}-r_{i}|\leq\delta$ and $|\widehat{b}_{i}-b_{i}|\leq\delta$ is enough to guarantee the sublinear regret stated in Theorem 3.1. In the subsection that follows, we show how to obtain such estimators in a bounded number of rounds.
@@ 3.2 Estimating the IV Rewards and ES
At a high level, it may seem impossible to disentangle learning the $r_{i}$’s and $b_{i}$’s just by observing the reward realization, which depends on their product. This hardness would disappear if one of the two (either the $r_{i}$’s or the $b_{i}$’s) were known. To see this, we present two simple thought experiments.
@@
Given Lemma 3.6 and Corollary 3.7, we now have a way to disentangle the learning of the $r_{i}$’s and the $b_{i}$’s. This is precisely the idea behind our algorithm: first, build estimators for the $r_{i}$’s and subsequently, use these when trying to infer the $b_{i}$’s. The tricky part arises because of the fact that the $r_{i}$’s and the $b_{i}$’s are connected multiplicatively. Note that for Lemmas 3.83.9 and 3.10 that follow, we use a fixed $\varepsilon$. We tune this $\varepsilon$ optimally in the end to obtain the no-regret guarantee.
Set $\varepsilon,\delta,M$ as stated in Theorem 3.1.
Initialize rounds $t=1$.
/* Build estimators $\{\widehat{r}_{i}\}_{i\in[K]}$/
@@
2       Initialize reward estimate $\widehat{r}_{i}=0$.
3       for blocks $j\in[M]$ do
@@
5                   Play arm $i_{R}$.
6                   Update $t\leftarrow t+1$.
            Play arm $i$, observe reward $R_{j}^{i}$, and update: $\widehat{r}_{i}\leftarrow\widehat{r}_{i}+\nicefrac{{R_{j}^{i}}}{{M}}$.
              // $q\approx 1-\varepsilon$
@@
9       Initialize state estimator $\widehat{v}_{i}=0$.
10       for pulls $1,\dots,N(\lambda)$ do
11             Play arm $i$.
12             Update $t\leftarrow t+1$.
13
@@
             Play arm $i$, observe reward $S_{j}^{i}$, and update: $\widehat{v}_{i}\leftarrow\widehat{v}_{i}+\nicefrac{{S_{j}^{i}}}{{M}}$.
              // Play $i$ when $q\approx b_{i}$
15
@@
17
Play arm $i_{R}$ for $N_{R}$ rounds, updating $t\leftarrow t+1$ after each one.
  // State $\geq 1-\varepsilon$
@@
The proofs of Lemmas 3.8 and 3.9 are based on an application of Hoeffding’s inequality combined with Corollary 3.7 to control the time it takes for the state to return to almost $1$.
@@
$\Pr[\mathcal{E}_{i}]\leq 2\exp\left(-2M\cdot(\delta-\varepsilon)^{2}\right)$ (4)
@@
$\Pr$ $\left[|e_{v}|\geq\frac{\delta}{2}\Big{|}|r_{i}|-|e_{r}|\Big{|}\Big{|}\mathcal{E}_{i}^{\prime}\right]\leq 2\exp\left(-2M\cdot\left(\frac{\delta}{2}\Big{|}|r_{i}|-|e_{r}|\Big{|}-\varepsilon\right)^{2}\right)$ (Lem. 3.9)
$=2\exp\left(-2M\cdot\left(\frac{\delta^{2}}{4}\Big{|}|r_{i}|-|e_{r}|\Big{|}^{2}-\delta\Big{|}|r_{i}|-|e_{r}|\Big{|}\varepsilon+\varepsilon^{2}\right)\right)$ ($||r_{i}|-|e_{r}||\leq 1$)
@@
where the last inequality is because $|e_{r}|\leq\delta$ (since we are conditioning on $\mathcal{E}_{i}^{\prime}$) and that $|r_{i}-\delta|\leq 1$ (since $r_{i}\in[0,1]$). From the law of total probability:
@@
Using Lemmas 3.8 and 3.10 we can prove Theorem 3.1 by bounding the number of rounds it takes for the estimators to converge to approximately correct values with high probability and the regret picked up in the event of failing to converge.
Sketch for the unknown $i_{R}$ case. Note that two instances of arms $\{(r_{i},b_{i})\}_{i\in[K]}$ and $\{(cr_{i},b_{i}/c)\}_{i\in[K]}$ for a scalar $c>0$ are equivalent. So, we can always scale $b_{i},\forall i\in[K]$ appropriately to make sure that we have a “replenishing” arm. The next part is to show how Algorithm 2 changes if we do not know which among the $K$ arms is the replenishing arm. The only thing that changes is the way that we estimate $\widehat{r}_{i}$’s; instead of using $i_{R}$ as the benchmark arm, we sample randomly an arm $z$. Then, after enough rounds, we can guarantee that with high probability $\hat{r}_{i}\to\bar{b}r_{i}$, where $\bar{b}=\sum_{j}\nicefrac{{b_{j}}}{{K}}$. The second part of Algorithm 2 remains the same. We can then guarantee that we have obtained estimates $\hat{b}_{i}\to\nicefrac{{b_{i}}}{{\bar{b}}}$ with high probability. Tuning again $\delta,\varepsilon,M$ we obtain the same order regret guarantee. The details and the new algorithm can be found in Appendix B.3.
@@ 4 Slow State Evolution: $\lambda\in[0,\widetilde{\Theta}(1/T)]$
For $\lambda=0$, the problem becomes an instance of the standard stochastic $K$-MAB, since the state is always $q_{t}=q_{0}=1,\forall t$. So, applying the standard UCB algorithm (Auer et al., 2002a) guarantees regret $R_{\textsc{DES}}(T)=\mathcal{O}(\sqrt{TK\log T})$. For the remainder of the section, we discuss the case where $\lambda\in(0,\widetilde{\Theta}(1/T)]$. Roughly, we show that applying EXP3.P (Auer et al., 2002b) “pretending” that there are no states (i.e., taking $\mathbb{E}[\tilde{r}_{i,t}]$ as being exogenously decided; see Algorithm 6 for formal description) incurs regret that is comparable to $R_{\textsc{DES}}(T)$ up to some factors that we formalize below.
@@
where $I^{*}$ is the best-fixed arm for the history of induced states being treated as exogenously given, i.e., $I^{*}=\arg\max_{i\in[K]}\sum_{t\in[T]}\tilde{r}_{i,t}=\arg\max_{i\in[K]}\sum_{t\in[T]}\tilde{q}_{t}^{i}\cdot r_{i}$, where $\tilde{q}_{t}^{i}$ is the value of the state if we play only arm $i$ from round $1$ to $t-1$. Using this, Equation (7) becomes:
$\sum_{t\in[T]}q_{t}^{i^{\star}}\cdot r_{i^{\star}}-\sum_{t\in[T]}\tilde{r}_{I_{t},t}\leq R_{\textsc{EXT}}(T)=\sum_{t\in[T]}q_{t}^{I^{\star}}\cdot r_{I^{\star}}-\sum_{t\in[T]}\tilde{r}_{I_{t},t}\leq\mathcal{O}\left(\sqrt{KT\log K}\right)$ (8)
@@
$R_{\textsc{DES}}(T)+\underbrace{\left[\sum_{t\in[T]}q_{t}^{i^{\star}}\cdot r_{i^{\star}}-\sum_{t\in[T]}q_{t}^{\pi^{\star}}r_{\pi^{\star}_{t}}\right]}_{A}\leq\mathcal{O}\left(\sqrt{KT\log K}\right)$
@@
$A$ $=\sum_{t\in[T]}q_{t}^{i^{\star}}\cdot r_{i^{\star}}-\sum_{t\in[T]}q_{t}^{\pi^{\star}}r_{\pi^{\star}_{t}}\geq\sum_{t\in[T]}(1-\lambda)^{t}r_{i^{\star}}-\sum_{t\in[T]}q_{t}^{\pi^{\star}}r_{\pi^{\star}_{t}}{}$ (9)
@@
where the first inequality is because $r_{i^{\star}}\geq r_{i},\forall i\in[K]$ and the second one because $1\geq\hat{q}_{t}$. Since $\{J_{t}\}_{t\in[T]}$ is any sequence, Equation (10) must also hold for the optimal sequence ${\pi^{\star}}$. In other words, $r_{i^{\star}}\geq q_{t}^{\pi^{\star}}r_{\pi^{\star}_{t}}$, $\forall t\in[T]$. Multiplying both sides with $(1-\lambda)^{t}$ and summing up for all $t\in[T]$, we get that $\sum_{t\in[T]}(1-\lambda)^{t}r_{i^{\star}}\geq\sum_{t\in[T]}(1-\lambda)^{t}q_{t}^{\pi^{\star}}r_{\pi^{\star}_{t}}$. Using the latter to relax the RHS of Equation (9) we get: